What is the pressure, in mmHg, of a 4.00 g sample of O2 gas, which has a temperature of 45.0 °C, and a volume of 3000 mL?
Answer #1

Molar mass of O2 = 32 g/mol

mass(O2)= 4.00 g

use:

number of mol of O2,

n = mass of O2/molar mass of O2

=(4 g)/(32 g/mol)

= 0.125 mol

Given:

V = 3000.0 mL

= (3000.0/1000) L

= 3 L

n = 0.125 mol

T = 45.0 oC

= (45.0+273) K

= 318 K

use:

P * V = n*R*T

P * 3 L = 0.125 mol* 0.08206 atm.L/mol.K * 318 K

P = 1.0873 atm

= 1.0873*760 mmHg

= 826 mmHg

Answer: 826 mmHg